If this person has a blood alcohol content of 0.082 % (w/v, or 0.082 g/100 mL), how many grams of alcohol would this person have in the blood? (Show work) [Hint: first calculate how many g of alcohol per liter, then how many g of alcohol in the total number of liters of blood].
1
Expert's answer
2019-09-17T04:55:36-0400

T=0.082 g/100ml = 0.00082 g/ml

If this person has 5 L (5000 ml) blood, we can determine mass of alcohol.

T= m/V m=T x V= 0.00082 x 5000 =4.1 g.